# Relaxed Equivariance via Multitask Learning

### v2

@@ Abstract
Incorporating equivariance as an inductive bias into deep learning architectures to take advantage of the data symmetry has been successful in multiple applications, such as chemistry and dynamical systems. In particular, roto-translations are crucial for effectively modeling geometric graphs and molecules, where understanding the 3D structures enhances generalization. However, equivariant models often pose challenges due to their high computational complexity. In this paper, we introduce REMUL, a training procedure for approximating equivariance with multitask learning. We show that unconstrained models (which do not build equivariance into the architecture) can learn approximate symmetries by minimizing an additional simple equivariance loss. By formulating equivariance as a new learning objective, we can control the level of approximate equivariance in the model. Our method achieves competitive performance compared to equivariant baselines while being 10×10$\times$10 × faster at inference and 2.5×2.5$\times$2.5 × at training.
@@ 1 Introduction
Equivariant machine learning models have achieved notable success across various domains, such as computer vision (Weiler et al., 2018; Yu et al., 2022), dynamical systems (Han et al., 2022; Xu et al., 2024), chemistry (Satorras et al., 2021; Brandstetter et al., 2022), and structural biology (Jumper et al., 2021). For example, incorporating equivariance w.r.t. translations and rotations ensures the correct handling of complex structures like graphs and molecules (Schütt et al., 2021; Bronstein et al., 2021; Thölke & Fabritiis, 2022; Liao et al., 2024). Equivariant machine learning models benefit from this inductive bias by explicitly leveraging symmetries of the data during the architecture design. Typically, such architectures have highly constrained layers with restrictions on the form and action of weight matrices and nonlinear activations (Batzner et al., 2022; Batatia et al., 2022). This may come at the expense of higher computational cost, making it sometimes challenging to scale equivariant architectures, particularly those relying on spherical harmonics and irreducible representations (Thomas et al., 2018; Fuchs et al., 2020; Liao & Smidt, 2023; Luo et al., 2024). On the other hand, equivariance constraints might limit the expressive power of the network, restricting its ability to act as a universal architecture (Dym & Maron, 2021; Joshi et al., 2023).
@@
Equivariant layers are not the only way to incorporate symmetries into deep neural networks. Several approaches have been proposed to either offload the equivariance restrictions to faster networks (Kaba et al., 2022; Mondal et al., 2023; Baker et al., 2024; Ma et al., 2024; Panigrahi & Mondal, 2024) or simplify the constraints by introducing averaging operations (Puny et al., 2022; Duval et al., 2023; Lin et al., 2024; Huang et al., 2024). Nonetheless, while these approaches leverage unconstrained architectures, they often require additional networks or averaging techniques to achieve equivariance and may not rely solely on adjustments to the training protocol. To this aim, a widely adopted strategy to replace ‘hard’ equivariance (i.e., built into the architecture itself) with a ‘soft’ one, is data augmentation (Quiroga et al., 2019; Bai et al., 2021; Gerken et al., 2022; Iglesias et al., 2023; Xu et al., 2023; Yang et al., 2024), whereby the training protocol of an arbitrary (unconstrained) network is augmented by assigning the same label to group orbits (e.g., rotated and translated versions of the input). In fact, recent works have shown that unconstrained architectures may offer a valid alternative provided that enough data are available (Wang et al., 2024; Abramson et al., 2024).
(a) Geometric Algebra Transformer
(b) Transformer
@@
Besides the challenges in computational cost and design, there are also tasks that do not exhibit full equivariance, such as dynamical phase transitions (Baek et al., 2017; Weidinger et al., 2017), polar fluids (Gibb et al., 2024), molecular nanocrystals (Yannouleas & Landman, 2000), and cellular symmetry breaking (Goehring et al., 2011; Mietke et al., 2019). For such tasks, fully-equivariant networks might be excessively constrained, which further motivates the design of a more flexible approach.
@@
In this work, we present REMUL: Relaxed Equivariance via Multitask Learning. REMUL is a training procedure that aims to learn approximate equivariance during training for unconstrained networks using a multitask approach with adaptive weights. We conduct a comprehensive evaluation of unconstrained models trained with REMUL, comparing their performance and computational efficiency against equivariant models. We consider Transformers and Graph Neural Networks (GNNs), as well as their E(3) equivariant versions, as our main baselines, focusing on roto-translation group.
Our contributions are as follows:
@@
  • We demonstrate that by adjusting the weighting of the equivariance loss, we can modulate the extent to which a model exhibits equivariance, depending on the requirements of the task. Specifically, tasks that demand full equivariance require a higher weight on the equivariance component, whereas tasks that require less strict equivariance can be managed with lower weights.
@@
  • By leveraging the efficiency of Transformers, we achieve up to 10×10$\times$10 × speed-up at inference and 2.5×2.5$\times$2.5 × speed-up in training compared to equivariant baselines. This finding could provide motivations for the use of unconstrained models, which do not require equivariance in their design, potentially offering a more practical approach.
  • We point out that the standard Transformer exhibits a more convex loss surface near the local minima compared to the Geometric Algebra Transformer (Brehmer et al., 2023), which can indicate further evidence of the optimization difficulties of equivariant networks.
@@ 2.1 Symmetry Groups and Equivariant Models
Definition 2.2.
A function $f$ is invariant w.r.t. the group $G$ if for all $g\in G$ and $x\in X$,
$f(\phi(g)(x))=f(x)$ (2)
@@ 2.2 Equivariance as a Constrained Optimization Problem
Consider a class of parametric functions $f_{\theta}$, typically implemented as neural networks, whose parameters $\theta$ are estimated via a general training objective based on data pairs $(x,y)\sim q$:
[…]
Here, $\mathcal{L}$ represents the loss function that quantifies the discrepancy between the model’s predictions $f_{\theta}(x)$ and the true labels $y$. The class of models is considered equivariant with respect to a group $G$ if it satisfies the constraint in Equation 1 for any input $x\in X$ and for any action $g\in G$.
Equivariance is typically achieved by design, by imposing constraints on the form of $f_{\theta}$. Since $f_{\theta}$ is usually composed of multiple layers, ensuring equivariance implies restrictions on the operations performed in each layer, a canonical example being message-passing graph neural networks whose local aggregations need to be permutation-equivariant to respect the overall invariance to the action of the symmetric group $S_{n}$. As such, finding an equivariant solution to the minimization problem in Equation LABEL:eq:_minimization_problem corresponds to solving the following constrained optimization:
[…]
@@
In general, such optimization is challenging, leading to complex design choices to enforce equivariance that could ultimately restrict the class of minimizers and make the training harder. Additionally, for relevant tasks, the optimal solution only needs to be approximate equivariant (Wang et al., 2022; Petrache & Trivedi, 2023; Kufel et al., 2024; Ashman et al., 2024) meaning that the extent to which a model needs to exhibit equivariance can vary significantly based on the specific characteristics of the data and the requirements of the downstream application. In light of these reasons, we necessitate a flexible approach to incorporating equivariance into the learning process. To address this, we propose REMUL, a training procedure that replaces the hard optimization problem with a soft constraint, by using a multitask learning approach with adaptive weights.
@@ 3.1 Equivariance as a New Learning Objective
Our main idea is to formulate equivariance as a multitask learning problem for an unconstrained model. We achieve that by relaxing the optimization problem in Equation 4. Namely, once we introduce a functional $\mathcal{F}_{\mathcal{X},G}$ that measures the equivariance of a candidate function $f_{\theta}$, we replace the constrained variational problem in Equation 4 with
[…]
@@
We note that in conventional supervised settings, one has access to a dataset $\mathcal{X}=\{x_{1},x_{2},\ldots,x_{n}\}$ with corresponding labels $\mathcal{Y}=\{y_{1},y_{2},\ldots,y_{n}\}$. We can then introduce
$\mathcal{L}_{\text{obj}}(f_{\theta},\mathcal{X},\mathcal{Y})=\sum_{i=1}^{n}\mathcal{L}(f_{\theta}(x_{i}),y_{i}),$ (6)
and formulate the optimization as:
@@
$\mathcal{L}_{\text{total}}(f_{\theta},\mathcal{X},\mathcal{Y},G)=\alpha\mathcal{L}_{\text{obj}}(f_{\theta},\mathcal{X},\mathcal{Y})+\beta\mathcal{L}_{\text{equi}}(f_{\theta},\mathcal{X},\mathcal{Y},G),$ (7)
@@
where $\mathcal{L}_{\text{equi}}(f_{\theta},\mathcal{X},\mathcal{Y},G)$ represents our augmented equivariance loss, specifically designed to enforce the model’s adherence to the symmetry action of the group $G$, given a dataset $\mathcal{X}$ and labels $\mathcal{Y}$. For a finite number of training samples $n$, we propose an equivariant loss $\mathcal{L}_{\text{equi}}$ of the form:
$\mathcal{L}_{\text{equi}}(f_{\theta},\mathcal{X},\mathcal{Y},G)=\sum_{i=1}^{n}\ell(f_{\theta}(\phi(g_{i})(x_{i})),\rho(g_{i})(y_{i}))$ (8)
@@
Here $\ell$ is a metric function, typically an $L_{1}$ or $L_{2}$ norm, that quantifies the discrepancy between $f(\phi(g_{i})(x_{i}))$ and $\rho(g_{i})(y_{i})$, with $g_{i}\in G$ randomly-selected group elements for each sample. In fact, in our implementation, we change the group elements being sampled in each training step.
@@
The parameters $\alpha$ and $\beta$ defined in Equation 7 are weighting factors that balance the traditional objective loss with the equivariance loss, enabling practitioners to tailor the training process according to specific requirements of symmetry and generalization. More specifically, a large value of $\beta$ indicates a more equivariant function while the smaller value of $\beta$ indicates a less equivariant function. These parameters allow us to control the trade-off between model generalization and equivariance, based on the specific requirements of the task.
@@ 3.2 Adapting Penalty Parameters during Training
For simultaneously learning the objective and equivariance losses, we consider two distinct approaches to regulate the penalty parameters $\alpha$ and $\beta$: constant penalty and gradual penalty. The constant penalty assigns a fixed weight to each task’s loss throughout the training process. In contrast, the gradual penalty dynamically adjusts the weights of each task’s loss during training. For gradual penalty, we use the GradNorm algorithm introduced by Chen et al. (2018), which is particularly suited for tasks that involve simultaneous optimization of multiple loss components, as it dynamically adjusts the weight of each loss during training. It updates the weights of the loss components based on the magnitudes of their gradients, w.r.t the last layer in the network, which is essential for the contribution of each loss. It also has a learning rate parameter $\eta$, that fine-tunes the speed at which the weights are updated, providing precise control over their convergence rates (see Algorithm 1 for details).
@@
1:Input: $\alpha$, $\beta$, $\eta$, $\gamma$, $\mathcal{L}_{\text{obj}}$, $\mathcal{L}_{\text{equi}}$, and $\mathcal{W}$ (the weights of the last layer in the network)
$\mathcal{G}_{\text{obj}}=\|\nabla_{\mathcal{W}}\alpha\mathcal{L}_{\text{obj}}\|_{2}$, $\tilde{\mathcal{L}_{\text{obj}}}=\mathcal{L}_{\text{obj}}/\mathcal{L}_{\text{obj}}(0)$
3:$\mathcal{G}_{\text{equi}}=\|\nabla_{\mathcal{W}}\beta\mathcal{L}_{\text{equi}}\|_{2}$, $\tilde{\mathcal{L}_{\text{equi}}}=\mathcal{L}_{\text{equi}}/\mathcal{L}_{\text{equi}}(0)$
4:$\bar{\mathcal{G}}=\frac{\mathcal{G}_{\text{obj}}+\mathcal{G}_{\text{equi}}}{2}$, $r=\frac{\tilde{\mathcal{L}_{\text{obj}}}+\tilde{\mathcal{L}_{\text{equi}}}}{2}$
5:$r_{\alpha}=\frac{\tilde{\mathcal{L}}_{\text{obj}}}{r}$, $r_{\beta}=\frac{\tilde{\mathcal{L}}_{\text{equi}}}{r}$
6:$\mathcal{L}_{\text{g}}=|\mathcal{G}_{\text{obj}}-\bar{\mathcal{G}}\times[r_{\alpha}]^{\gamma}|+|\mathcal{G}_{\text{equi}}-\bar{\mathcal{G}}\times[r_{\beta}]^{\gamma}|$
7:$\alpha=\alpha-\eta\nabla_{\alpha}\mathcal{L}_{\text{g}}$
8:$\beta=\beta-\eta\nabla_{\beta}\mathcal{L}_{\text{g}}$
9:Return: $\alpha$, $\beta$
3.3 Equivariance with Data Augmentation
Data augmentation is a widely recognized technique that enhances the performance of machine learning models by including different transformations in the training process. It involved creating a transformed input and measuring the original loss between the model prediction and the transformed target. In contrast, our method utilizes an additional controlled equivariance loss to incorporate symmetrical considerations simultaneously with the objective loss during training. In fact, traditional data augmentation techniques can be interpreted as special cases of Equation 7 where $\alpha=0$ and $\beta=1$.
4 Quantifying Learned Equivariance
Using group transformations to measure and assess the symmetries of ML models has been studied in several domains (Lyle et al., 2020; Kvinge et al., 2022; Moskalev et al., 2023; Gruver et al., 2023; Speicher et al., 2024). Inspired by the idea of frame-averaging (Puny et al., 2022; Duval et al., 2023; Lin et al., 2024), in this section, we introduce a metric to quantify the degree of equivariance exhibited by a function $f$.
Starting from Equation 1, the group integration of both sides w.r.t. the normalized Haar measure $\mu$ yields:
[…]
@@ 4 Quantifying Learned Equivariance
Where $M$ is a large number of samples from $G$. Given the group averages, we define the equivariance error $E(f,G)$ as the average norm of the difference between these two averages over the data distribution $D$:
$E(f,G)=\frac{1}{|D|}\sum_{x\in D}\left\|\frac{1}{M}\sum_{i=1}^{M}\rho(g_{i})(f(x))-\frac{1}{M}\sum_{i=1}^{M}f(\phi(g_{i})(x))\right\|_{2}$ (12)
Here $\|\cdot\|_{2}$ denotes an $L_{2}$ norm (for non-scalar function). This error indicates the average deviation of a function $f$ from perfect equivariance across the data distribution $D$ (lower value means more equivariant function).
We also propose another measure that takes the average over the group of differences between $f(\phi(g)(x))$ and $\rho(g)(f(x))$,
$E^{\prime}(f,G)=\frac{1}{|D|}\sum_{x\in D}\frac{1}{M}\sum_{i=1}^{M}\left\|f(\phi(g_{i})(x))-\rho(g_{i})(f(x))\right\|_{2}$ (13)
Equation 12 & Equation 13 indicate a practical metric for evaluating how closely the function $f$ approximates perfect equivariance throughout a data distribution $D$ (which should be zero for a perfect equivariance function). In practice, we use $M=100$ samples from the group and noticed this was sufficient to obtain stable results. We also observed that both measures have very similar behavior in our experiments, where $E$ and $E^{\prime}$ are near zero for equivariant models. We also demonstrate that increasing the value of $\beta$ in Equation 7 results in a less equivariant error for $E$ and $E^{\prime}$.

### v3

@@ Abstract
[…]
@@ 1 Introduction
Equivariant machine learning models have achieved notable success across various domains, such as computer vision [1, 2], dynamical systems [3, 4], chemistry [5, 6], and structural biology [7]. For example, incorporating equivariance w.r.t. translations and rotations ensures the correct handling of complex structures like graphs and molecules [8, 9, 10, 11]. Equivariant machine learning models benefit from this inductive bias by explicitly leveraging symmetries of the data during the architecture design. Typically, such architectures have highly constrained layers with restrictions on the form and action of weight matrices and nonlinear activations [12, 13]. This may come at the expense of higher computational cost, making it sometimes challenging to scale equivariant architectures, particularly those relying on spherical harmonics and irreducible representations [14, 15, 16, 17]. On the other hand, equivariance constraints might limit the expressive power of the network, restricting its ability to act as a universal architecture [18].
@@
Equivariant layers are not the only way to incorporate symmetries into deep neural networks. Several approaches have been proposed to either offload the equivariance restrictions to faster networks [19, 20, 21, 22, 23] or simplify the constraints by introducing averaging operations [24, 25, 26, 27]. Nonetheless, while these approaches leverage unconstrained architectures, they often require additional networks or averaging techniques to achieve equivariance and may not rely solely on adjustments to the training protocol. To this aim, a widely adopted strategy to replace ‘hard’ equivariance (i.e., built into the architecture itself) with a ‘soft’ one, is data augmentation [28, 29, 30, 31, 32, 33, 34, 35, 36], whereby the training protocol of an arbitrary (unconstrained) network is augmented by assigning the same label to group orbits (e.g., rotated and translated versions of the input). In fact, recent works have shown that unconstrained architectures may offer a valid alternative, provided that enough data are available [37, 38].
@@
Besides the challenges in computational cost and design, there are also tasks (especially in scientific applications of ML) that do not exhibit full equivariance, such as dynamical phase transitions [39, 40], polar fluids [41], molecular nanocrystals [42], and cellular symmetry breaking [43, 44]. For such tasks, fully-equivariant networks might be excessively constrained, which further motivates the design of a more flexible approach.
@@
In this work, we present REMUL: Relaxed Equivariance via Multitask Learning. REMUL is a training procedure that aims to learn approximate equivariance during training for unconstrained networks using a multitask approach with adaptive weights. We conduct a comprehensive evaluation of unconstrained models trained with REMUL, comparing their performance and computational efficiency to equivariant models. We consider Transformers and Graph Neural Networks (GNNs) and their roto-translational (E(3))-equivariant versions as our main baselines. Our contributions are:
@@
  • We demonstrate that by adjusting the weighting of the equivariance loss, we can modulate the extent to which a model exhibits equivariance, depending on the task’s requirements. Specifically, tasks that demand full equivariance require a higher weight on the equivariance term, whereas tasks that require less strict equivariance can be managed with lower weights.
@@
  • By leveraging the efficiency of Transformers, we achieve up to $10\times$ speed-up at inference and $2.5\times$ speed-up in training compared to equivariant baselines. This finding could provide motivations for the use of unconstrained models, which do not require equivariance in their design, potentially offering a more practical approach.
  • We point out that the standard Transformer exhibits a more convex loss surface near the local minima compared to the Geometric Algebra Transformer [45], which can indicate further evidence of the optimization difficulties of equivariant networks.
@@ 2.1 Symmetry Groups and Equivariant Models
Definition 2.
A function $f$ is invariant w.r.t. the group $G$ if for all $g\in G$, $x\in X$$f(\phi(g)(x))=f(x).$
@@ 2.2 Equivariance as a Constrained Optimization Problem
Consider a class of parametric functions $f_{\theta}$ from a hypothesis space $\mathcal{H}$, typically implemented as neural networks, whose parameters $\theta$ are estimated via a general training objective based on data pairs $(x,y)\sim q$:
$\underset{\theta}{\text{minimize}}$ $\mathbb{E}_{(x,y)\sim q}\left[\mathcal{L}(f_{\theta}(x),y)\right]$ (2)
Here, $\mathcal{L}$ represents the loss function that quantifies the discrepancy between the model’s predictions $f_{\theta}(x)$ and the true labels $y$. The class of models is considered equivariant with respect to a group $G$ if it satisfies the constraint in Eq. 1 for any input $x\in X$ and for any action $g\in G$.
Equivariance is typically achieved by design, by imposing constraints on the form of $f_{\theta}$. Since $f_{\theta}$ is usually composed of multiple layers, ensuring equivariance implies restrictions on the operations performed in each layer, a canonical example being message-passing graph neural networks whose local aggregations need to be permutation-equivariant to respect the overall invariance to the action of the symmetric group $S_{n}$. As such, finding an equivariant solution to the minimization problem in Eq. LABEL:eq:_minimization_problem corresponds to solving the following constrained optimization:
$\underset{\theta}{\text{minimize}}$ $\mathbb{E}_{(x,y)\sim q}\left[\mathcal{L}(f_{\theta}(x),y)\right]$ (3)
@@
In general, such optimization is challenging, leading to complex design choices to enforce equivariance that could ultimately restrict the class of minimizers and make the training harder. Additionally, for relevant tasks, the optimal solution only needs to be approximate equivariant [46, 47, 48, 49] meaning that the extent to which a model needs to exhibit equivariance can vary significantly based on the specific characteristics of the data and the requirements of the downstream application. In light of these reasons, we require a flexible approach to incorporating equivariance into the learning process. To address this, we propose REMUL, a training procedure that replaces the hard optimization problem with a soft constraint, by using a multitask learning approach with adaptive weights.
@@ 3.1 Equivariance as a New Learning Objective
Our main idea is to formulate equivariance as a multitask learning problem for an unconstrained model. We achieve that by relaxing the optimization problem in Eq. 3. Namely, once we introduce a functional $\mathcal{F}_{\mathcal{X},G}$ that measures the equivariance of a candidate function $f_{\theta}$, we replace the constrained variational problem in Eq. 3 with
$\underset{\theta}{\text{minimize}}\,\,\mathbb{E}_{(x,y)\sim q}\left[\alpha\mathcal{L}(f_{\theta}(x),y)+\beta\mathcal{F}_{\mathcal{X},G}(f_{\theta}(x),y)\right],$ (4)
@@
Empirical Formulation. Let $\mathcal{D}_{n}=\{(x_{i},y_{i})\}_{i=1}^{n}$ be a training sample of size $n$ drawn i.i.d. from an underlying distribution $P_{XY}$ on $\mathcal{X}\times\mathcal{Y}$. In conventional supervised settings, we define the empirical version of our optimization problem as:
@@
$\mathcal{L}_{\text{total}}(f_{\theta},\mathcal{X},\mathcal{Y},G)=\alpha\widehat{\mathcal{L}}_{\mathrm{obj}}(f_{\theta},\mathcal{X},\mathcal{Y})+\beta\widehat{\mathcal{L}}_{\mathrm{equi}}(f_{\theta},\mathcal{X},\mathcal{Y},G),$ (5)
@@
where $\widehat{\mathcal{L}}_{\mathrm{obj}}(f_{\theta},\mathcal{X},\mathcal{Y})$ is the empirical objective loss given by $\widehat{\mathcal{L}}_{\mathrm{obj}}(f_{\theta},\mathcal{X},\mathcal{Y})=\tfrac{1}{n}\sum_{i=1}^{n}\mathcal{L}(f_{\theta}(x_{i}),y_{i})$, and $\widehat{\mathcal{L}}_{\mathrm{equi}}(f_{\theta},\mathcal{X},\mathcal{Y},G)$ represents our augmented equivariance loss, specifically designed to enforce the model’s adherence to the symmetry action of the group $G$. For a finite number of training samples $n$, we propose an empirical equivariant loss $\widehat{\mathcal{L}}_{\mathrm{equi}}$ of the form:
$\widehat{\mathcal{L}}_{\mathrm{equi}}(f_{\theta},\mathcal{X},\mathcal{Y},G)=\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{g\sim G}\Bigl[\ell(f_{\theta}(\phi(g)x_{i}),\;\rho(g)y_{i})\Bigr]$ (6)
@@
here $\ell$ is a metric function, typically an $L_{1}$ or $L_{2}$ norm, that quantifies the discrepancy between $f(\phi(g)(x_{i}))$ and $\rho(g)(y_{i})$, with $g\in G$ randomly-selected group elements drawn from a uniform distribution for each sample. In our implementation, we enhance computational efficiency by selecting a single group element per sample at each training step, which we found produces effective results. In addition, we show how the performance varies as we increase the number of group samples.
@@
Characterizing the REMUL Trade-off. While REMUL is presented as a practical training procedure, it can be theoretically understood as a regularized optimization problem. The parameters $\alpha$ and $\beta$ defined in Eq. 5 are weighting factors that balance the traditional objective loss with the equivariance loss, enabling practitioners to tailor the training process according to specific requirements of symmetry and generalization. The following proposition characterizes the properties of the empirical minimizer $f_{\alpha,\beta}$ and the underlying trade-offs between task performance and equivariance. The proof is provided in Appendix B.
Proposition 1.
Let $f_{\alpha,\beta}\in\arg\min_{f\in\mathcal{H}}\widehat{\mathcal{L}}_{\mathrm{total}}(f;\alpha,\beta)$ be an empirical minimizer of the REMUL objective, and let $f^{\star}_{\mathrm{obj}}\in\arg\min_{f\in\mathcal{H}}\widehat{\mathcal{L}}_{\mathrm{obj}}(f)$ be an empirical minimizer for the objective loss alone. Then:
  1. (a)
    $f_{\alpha,\beta}$ is Pareto optimal for the bi-objective problem $(\min\widehat{\mathcal{L}}_{\mathrm{obj}}(f),\min\widehat{\mathcal{L}}_{\mathrm{equi}}(f))$.
  2. (b)
    The following trade-off inequality holds:
    $\widehat{\mathcal{L}}_{\mathrm{obj}}(f_{\alpha,\beta})-\widehat{\mathcal{L}}_{\mathrm{obj}}(f^{\star}_{\mathrm{obj}})\leq\frac{\beta}{\alpha}\left(\widehat{\mathcal{L}}_{\mathrm{equi}}(f^{\star}_{\mathrm{obj}})-\widehat{\mathcal{L}}_{\mathrm{equi}}(f_{\alpha,\beta})\right).$ (7)
Controlling Approximate Equivariance via $\beta/\alpha$.
Eq. 7 quantifies the empirical cost of enforcing equivariance, showing that any increase in primary task’s loss beyond the unconstrained minimum $\widehat{\mathcal{L}}_{\mathrm{obj}}(f^{\star}_{\mathrm{obj}})$) is bounded by the product of relative weight $\beta/\alpha$ and the achieved reduction in the equivariance loss (from $\widehat{\mathcal{L}}_{\mathrm{equi}}(f^{\star}_{\mathrm{obj}})$ down to $\widehat{\mathcal{L}}_{\mathrm{equi}}(f_{\alpha,\beta})$). The ratio $\beta/\alpha$ serves as a lever to control the solution’s properties: When $\beta/\alpha\to 0$, the objective prioritizes task performance, causing $f_{\alpha,\beta}$ to approximate $f^{\star}_{\mathrm{obj}}$ (potentially sacrificing equivariance if $f^{\star}_{\mathrm{obj}}$ lacks natural symmetry). In contrast, when $\beta/\alpha\to\infty$, the objective prioritizes equivariance, driving $\widehat{\mathcal{L}}_{\mathrm{equi}}(f_{\alpha,\beta})$ toward zero (at the cost of task performance). Finally, at intermediate $\beta/\alpha$ values, the solution $f_{\alpha,\beta}$ represents a specific balance on the empirical Pareto frontier. REMUL thus allows learning a tunable degree of approximate equivariance, with larger $\beta$ producing a more equivariant function, while smaller $\beta$ produces a less equivariant function. This flexibility allows us to control the trade-off between model generalization & equivariance based on the task’s requirements, as we demonstrate empirically in Section 6.
@@ 3.2 Adapting Penalty Parameters during Training
For simultaneously learning the objective and equivariance losses, we consider two distinct approaches to regulate the penalty parameters $\alpha$ and $\beta$: constant penalty and gradual penalty. The constant penalty assigns a fixed weight to each task’s loss throughout the training process. In contrast, the gradual penalty dynamically adjusts the weights of each task’s loss during training. For gradual penalty, we use the GradNorm algorithm introduced by [50], which is particularly suited for tasks that involve simultaneous optimization of multiple loss components, as it dynamically adjusts the weight of each loss during training. It updates the weights of the loss components based on the magnitudes of their gradients, w.r.t the last layer in the network, which is essential for the contribution of each loss. It also has a learning rate parameter $\eta$, that fine-tunes the speed at
@@
1:Input: $\alpha$, $\beta$, $\eta$, $\gamma$, $\widehat{\mathcal{L}}_{\mathrm{obj}}$, $\widehat{\mathcal{L}}_{\mathrm{equi}}$, and $\mathcal{W}$ (the weights of the last layer in the network)
@@
[…]
9:Return: $\alpha$, $\beta$
which the weights are updated, providing precise control over their convergence rates (see Algorithm 1 for details).
Equivariance with Data Augmentation.
Standard data augmentation for enforcing equivariance typically involves augmenting the training data with pairs $(\phi(g)(x_{i}),\rho(g)(y_{i}))$ and training the model $f_{\theta}$ using only the original task loss $\mathcal{L}_{obj}$, i.e., minimizing $\sum_{i}\mathcal{L}(f_{\theta}(\phi(g)(x_{i})),\rho(g)(y_{i}))$ over the augmented dataset. This implicitly encourages the network to learn symmetries by penalizing predictions on transformed data using the standard task objective. REMUL differs by introducing a separate, explicit equivariance loss term $\mathcal{L}_{\text{equi}}$ alongside the standard objective loss $\mathcal{L}_{\text{obj}}$ on the original data, as indicated in Eq. 5. The multitask framework with weights $\alpha,\beta$ allows explicit control over the balance between fitting the original data and enforcing the equivariance constraint.
4 Quantifying Learned Equivariance
Using group transformations to measure and assess the symmetries of ML models has been studied in several domains [51, 52, 53, 54, 55]. Inspired by the idea of frame-averaging [24, 25, 26], we introduce a metric to quantify the degree of equivariance exhibited by a function $f$, defined as:
$E(f,G)=\frac{1}{|\mathcal{D}|}\sum_{x\in\mathcal{D}}\left\|\frac{1}{M}\sum_{i=1}^{M}\rho(g_{i})(f(x))-\frac{1}{M}\sum_{i=1}^{M}f(\phi(g_{i})(x))\right\|_{2}$ (8)
where $\|\cdot\|_{2}$ denotes an $L_{2}$ norm (for non-scalar function), and $M$ is a large number of samples from $G$. (Proof in Appendix B). This error indicates the average deviation of a function $f$ from perfect equivariance across the data distribution $\mathcal{D}$ (lower value means more equivariant function). We also compare to the standard measure that takes the average over the group of differences between $f(\phi(g)(x))$ and $\rho(g)(f(x))$,
$E^{\prime}(f,G)=\frac{1}{|\mathcal{D}|}\sum_{x\in\mathcal{D}}\frac{1}{M}\sum_{i=1}^{M}\left\|f(\phi(g_{i})(x))-\rho(g_{i})(f(x))\right\|_{2}$ (9)
We observed that both measures have very similar behavior in our experiments, where $E$ and $E^{\prime}$ are near zero for equivariant models. Furthermore, as we discussed in Section 3, we demonstrate empirically that increasing the REMUL penalty weight $\beta$ (Eq. 5) results in a lower equivariant error for $E$ and $E^{\prime}$.